# A Baseline Method for Removing Invisible Image Watermarks using Deep Image Prior

### v1

@@ 5 Conclusion and discussion
Acknowledgments
Liang H. is partially supported by NIH fund R01NS131314. Sun J. is partially supported by NIH funds R01NS131314 and R01CA287413. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. The authors acknowledge the Minnesota Supercomputing Institute (MSI) at the University of Minnesota for providing resources that contributed to the research results reported in this article.
@@ Appendix B Ranges of hyperparameters of various evasion methods
In Table 5, the quality index of VAE indexes the individual pretrained model (six provided in the original work). Thus, the minimal resolution for VAE search is $1$ and cannot be smaller. Moreover, the last column shows the typical value used in the existing literature for robustness evaluation as a reference, e.g., Zhao et al. (2023); Saberi et al. (2023). To the best of our knowledge, the reason why these values are chosen as defaults is not well justified.
Appendix C Image quality v.s. evasion success
@@
Appendix D Additional visualization of evasion image quality on rivaGAN

### v2

@@ 5 Conclusion and discussion
6 Ethical statement
One potential ethical concern regarding this paper is that it could facilitate the unauthorized removal of watermarks, thereby enabling copyright infringement. However, this concern may be unfounded: Tutorials on using Deep Image Prior (DIP) to remove visible watermarks (e.g., Rishik (2020)) are already publicly available; While research on invisible watermarking is active, such methods have not yet been widely deployed in real-world applications. Rather than promoting misuse, our work aims to proactively identify limitations in current invisible watermarking techniques and to offer concrete recommendations for developing more robust watermarking strategies.
@@ Appendix A DALLE-2 visible watermark
An example of a visible DALLE-2 watermark in colored blocks is shown Fig. 9.
@@ Appendix B Ranges of hyperparameters of various evasion methods
In Table 6, the “quality index” for the VAE refers to the selection of a specific pretrained model (among the six provided in the original work), hence the minimal grid resolution for this hyperparameter is 1. The final column lists the default values commonly used in the literature for robustness evaluation (e.g., Zhao et al. (2024b); Saberi et al. (2023)). To the best of our knowledge, the rationale behind the choice of these default settings is not well documented.
Appendix C Runtime comparison of different watermark evasions
Table 7: Runtime comparison of evading a single watermarked image by different watermark evasion methods
Method Evasion configuration Time (s)
Diffuser 10 different diffusion steps 3.85
BM3D 5 different standard deviation parameters 18.2
DIP 500 iterations 26.6
DiffPure 10 different diffusion steps 290.75
We perform a runtime comparison of different watermark evasion methods mentioned in this paper. The experiment is carried out on a local desktop with Windows OS equipped with an Intel Core i7-12700K processor and an NVIDIA RTX 3080 GPU, on the evasions of a single image. The result can be found in Table 7: Our DIP-based evasion is comparable to Diffuser and BM3D, and is substantially faster than DiffPure. Moreover, we note that recent research has tried to accelerate DIP-based denoising, e.g., Li et al. (2023a), which—although beyond the scope of this paper—represents promising directions for further improving the computational efficiency of our DIP-based watermark evasion.
Appendix D Image quality v.s. evasion success
@@ Appendix C Runtime comparison of different watermark evasions
Fig. 10 presents several illustrative trajectories of the relationship between image quality and the evasion success evaluated on our DIP-based evasion.
Appendix E Additional visualization of evasion image quality on rivaGAN
@@
Fig. 11 presents evasion images with the best visual quality from the different evasion methods considered in this paper. Note that all other methods except for DIP results in visual artifacts in the evasion image, e.g., pixel jitters (Gaussian Noise, JPEG and Diffuser) or overly smoothing effect (bm3d, DiffPure and VAE).
Appendix F Additional discussion on the WIND watermark
Arabi et al. (2024) proposes an inpainting-based variant of the WIND watermark ($\text{WIND}_{\text{inpainting}}$), whose underlying principle is similar to that of TreeRing (Wen et al., 2023) and aims to extend diffusion-based in-processing watermarking methods to post-processing scenarios. The primary focus of Arabi et al. (2024) is the identification (i.e., recovery) of the watermark key, but not watermark detection—the binary classification of watermarked vs. non-watermarked images; see the evaluation in Table 1 in Wen et al. (2023).
To evaluate the detection performance of $\text{WIND}_{\text{inpainting}}$, we perform the following experiment: We sample 100 images from the COCO dataset and generate the corresponding watermarked versions using $\text{WIND}_{\text{inpainting}}$ (Arabi et al., 2024). We then apply the $\text{WIND}_{\text{inpainting}}$ decoder to all 200 images (clean and watermarked) and record the first-stage distance for each. As described in Arabi et al. (2024), this distance serves as the basis for watermark detection. We plot the histogram of the first-stage distance for clean and watermarked images, respectively, as shown in Fig. 12.
Figure 12: Distribution of the first-stage distance for clean and watermarked images decoded by $\text{WIND}_{\text{inpainting}}$
We observe that thresholding the first-stage distance produced by the $\text{WIND}_{\text{inpainting}}$ decoder is insufficient for reliably distinguishing between clean and watermarked images. Consequently, we do not include $\text{WIND}_{\text{inpainting}}$ in our comparison.